OpenStudy (anonymous):

cos(285[degrees]) Exact value?

OpenStudy (anonymous):

cos 285 = 0.25882

OpenStudy (anonymous):

cos(285) =cos(360-285) =cos(75) =cos(45+30): Use rule for cos(theta +phi) and get exact value.